Explain if the limit of the function f(t) exist at t=1 where,

F(t)={ t2+5 if t<1} 

 {3-3t if t>1}



1
Expert's answer
2021-02-24T07:14:59-0500

The limit of "f(t)" at "t=1" exists if the one-sided limits exist at "t=1" , and are equal.

For given function

"f(t)=\left\{\begin{array}{ll}t^2+5,&t<1,\\
3-3t,&t>1\end{array}\right."

the left-hand limit equals

"\displaystyle\lim\limits_{t\to1^-}f(t)=
\lim\limits_{t\to1}(t^2+5)=1^2+5=1+5=6",

the right-hand limit equals

"\displaystyle\lim\limits_{t\to1^+}f(t)=
\lim\limits_{t\to1}(3-3t)=3-3\cdot1=3-3=0".

So, the one-sided limits of given function exist at "t=1" , but are unequal

"\displaystyle\lim\limits_{t\to1^-}f(t)\neq\lim\limits_{t\to1^+}f(t)".

Hence, the limit of the function at "t=1" does not exist.